Calculate Payroll K. Mello Company has three employees-a consultant, a computer programmer, and an administrator. The following payroll information is available for each employee: Consultant Computer Programmer Administrator Regular earnings rate $2,710 per week $34 per hour $50 per hour Overtime earnings rate Not applicable 2 times hourly rate 1.5 times hourly rate Federal income tax withheld $910 $252 $500 For hourly employees, overtime is paid for hours worked in excess of 40 hours per week. For the current pay period, the computer programmer worked 53 hours and the administrator worked 63 hours. Assume further that the social security tax rate was 6%, and the Medicare tax rate was 1.5%. Determine the gross pay and the net pay for each of the three employees for the current pay period. Assume the normal working hours in a week are 40 hours. If required, round your answers to two decimal places.
Answer #1

1. Consultant Income

Consultant's earning = $2710

Assumption - Consultant has worked for 40 hours as the same is not mentioned in the Question and Overtime is also not applicable as given in question.

Withheld tax = $910

Social security tax = 6% of Total earning = $2710 * 6% = $162.60

Medicare tax = 1.5% of total earning = $2710 * 1.5% = $40.65

Gross Pay = $2710

Net Pay = Gross pay - Withheld tax - Medicare tax - social security tax

Net Pay = $2710 - $910 - $40.65 - $162.60 = $1596.75

2. Computer Programmer

Rate of earning = $34 per Hour

Working Hours = 53 Hours

Computer programmer Earning = $34 * 53 = $1802

Overtime hours = 53 - 40 = 13 hours

Overtime earning = 2 times of hourly rate = 13 * $34 = $442 (once already added above)

Withheld tax = $252

Social security tax = $ 2244 * 6% = $134.64

Medicare tax = 1.5% of total earning = $2244 * 1.5% = $33.66

Gross Pay would be = $2244

Net Pay = Gross pay - Withheld tax - Medicare tax - social security tax

Net Pay = $2244 - $252 - $33.66 - 134.64 = $1823.70

3. Administrator Income

Rate of earning = $50 per Hour

Working Hours = 63 Hours

Computer programmer Earning = $50 * 63 = $3150

Overtime hours = 63 - 40 = 23 hours

Overtime earning = 1.5 times of hourly rate = 23 * $25 = $575 (once already added above so half of the rate is taken)

Withheld tax = $500

Social security tax = $ 3725 * 6% = $223.50

Medicare tax = 1.5% of total earning = $3725 * 1.5% = $55.88

Gross Pay would be = $3725

Net Pay = Gross pay - Withheld tax - Medicare tax - social security tax

Net Pay = $3725 - $500 - $55.88 - $223.50 = $2945.62

Thanks and Regards